A solid cylinder has a density of 8.9g/cm^3 and weighs 72g. Find its apparent weight when immersed totally in a liquid of density 0.8g/cm^3.
1
Expert's answer
2020-11-18T14:52:13-0500

By the definition of the apparent weight, we have:


"W_{A}=W_{c}-F_B,"

here, "W_{A}" is the apparent weight of the solid cylinder, "W_c" is the weight of the cylinder in air, "F_B=\rho_{liquid}V_{liquid}g" is the buoyant force that acts on the solid cylinder when it submerged in liquid, "\rho_{liquid}=800\ \dfrac{kg}{m^3}" is the density of the liquid, "V_{liquid}=V_{c}" is the volume of the liquid displaced that is equal to the volume of the solid cylinder and "g=9.8\ \dfrac{m}{s^2}" is the acceleration due to gravity.

Let's first find the volume of the liquid displaced:


"V_{liquid}=V_{c}=\dfrac{m}{\rho},""V_{liquid}=\dfrac{0.072\ kg}{8900\ \dfrac{kg}{m^3}}=8.1\cdot 10^{-6}\ m^3."

Finally, we can find the apparent weight of the solid cylinder:


"W_{A}=W_{c}-F_B,""W_{A}=mg-\rho_{liquid}V_{liquid}g,"

"W_{A}=0.072\ kg\cdot 9.8\ \dfrac{m}{s^2}-800\ \dfrac{kg}{m^3}\cdot 8.1\cdot 10^{-6}\ m^3\cdot 9.8\ \dfrac{m}{s^2}=0.64\ N."

Answer:

"W_{A}=0.64\ N."